How long will it take for your money to double if it were invested at 60% interest, compunded continuously?
Answer by stanbon(75887)

How long will it take for your money to double if it were invested at 60% interest, compunded continuously?
-------------
A(t) = Pe^(rt)
If it doubles A(t) = 2P
2P = Pe^(rt)
2 = e^(0.60t)
Take the natural log to get:
ln2 = 0.60t
t = [ln2]/0.60
t = 1.155 years
==============
Cheers,
Stan H.